OpenStudy (anonymous):

A right triangle has a hypotenuse of length 39 and a leg of length 15. What is the length of the other leg? If necessary, round your answer to two decimal places.

OpenStudy (ddcamp):

(leg)^2 + (leg)^2 = (hypotenuse)^2 a^2 + b^2 = c^2